# Random Projection Layers for Multidimensional Time Series Forecasting

### v2

@@ 5.2 Baseline Method
5.3 Benchmark Result
The benchmark results are summarized in Table 1. Our discussion on the performance differences among various methods will proceed in three parts: firstly, we will discuss the 11 baselines included in the original LargeST benchmark Liu et al. (2023); next, we will address the three new time series forecasting baselines we added to the benchmark; finally, we will highlight the performance of our proposed RPMixer method in comparison to the baselines.
[…]
@@ 5.4 Ablation Study
Figure 7: The ablation study result measured using average MAE, RMSE, and MAPE. Removing each component resulted in a degradation of the performance.
The pre-activation design emerged as the most consequential design decision. Upon its replacement with a post-activation design, we effectively removed the most important element of the original design, leading to a significant performance decrease. This decline is likely due to the elimination of identity mapping connections from the network. These connections are vital for the ensemble interpretation of the residual connection. Without them, the inclusion of a random projection layer no longer makes sense, as it weakens the ensemble-like-like behavior of the residual blocks (i.e., the mixer blocks in RPMixer), which was the motivation behind the design of the random projection layer.
The random projection layer design was the second most important decision in terms of performance difference. We disabled the random projection layer by converting it into a regular linear layer with trainable weights. This layer helps each mixer block in focusing on different aspects of the inter-node relationship, so its importance is not surprising. To verify our claim that the random projection layer promotes diversity, we constructed a correlation-error diagram comparing the proposed method with the variant without random projection layer. This diagram serves a similar purpose to the kappa-error diagram Margineantu and Dietterich (1997) used for analyzing base learners in an ensemble model. Before discussing the diagram, let us first introduce what a correlation-error diagram is.
Figure 8: The correlation-error diagram illustrates the interplay of diversity and error (MAE) across individual blocks in the model.
A correlation-error diagram is a visualization tool akin to the kappa-error diagram Margineantu and Dietterich (1997), used for analyzing the trade-off between performance and diversity in ensemble models. Each dot in the diagram represents a pair of base learners, with the $y$-axis showing the average performance of the pair’s outputs and the $x$-axis indicating the degree of agreement between them. Specifically, we use the Pearson correlation coefficient to measure agreement between the outputs of a pair of mixer blocks (i.e., the base learners). We use MAE, RMSE, and MAPE as performance measures. The figure created using MAE and Pearson correlation coefficient is shown in Fig. 8. Figures using other performance measures are included in Appendix A.5. We observe that the dots associated with the proposed method with a random projection layer have a wider distribution (larger range in $y$-axis) compared to the variant without the random projection layer. This confirms that our random projection layer effectively increases the diversity of the intermediate representation.
@@
5.5 Parameter Sensitivity Analysis
In our parameter sensitivity analysis, we focus on two hyper-parameters: the number of mixer blocks and the number of neurons in the random projection layer. We report performance on both the validation and test sets to demonstrate the generalizability of our findings across different data partitions. Fig. 8(a) illustrates the parameter sensitivity analysis for the number of mixer blocks.
(a) Setting the number of block to eight generally yields better results. This observation holds true for both validation and test data.
(b) The number is set by $m_{neuron}\sqrt{n}$, where $n$ is the number of nodes in the graph and $m_{neuron}$ is a factor. Setting this factor to one often has the best performance. This observation holds true for both validation and test data.
Figure 9: The parameter sensitivity analysis for the number of block and number of neurons parameter measured using average MAE, RMSE, and MAPE.
@@ 5.5 Parameter Sensitivity Analysis
For the number of neurons in the random projection layer, we set it as a function with respect to the number of nodes in the spatial-temporal data, as more neurons may be required for datasets with larger number of nodes. If the graph has $n$ nodes, we set the number of neurons in the random projection layer as $m_{neuron}\sqrt{n}$, where $m_{neuron}$ is the hyper-parameter controlling the number of neurons in the layer. This design allows the model to automatically use more neurons for datasets with more number of nodes, even if the hyper-parameter $m_{neuron}$ is set to the same value for all datasets. Fig. 8(b) displays the results of the parameter sensitivity analysis for the hyper-parameter $m_{neuron}$.
@@
6 Conclusion
In this paper, we proposed RPMixer, an all-MLP mixer model that incorporates random projection layers. These layers enhance the diversity among the outputs of each mixer block, thereby improving the overall performance of the model. Our experiments demonstrated that the random projection layers not only improve the diversity of the intermediate representation but also boost the model’s overall performance on large-scale spatial-temporal benchmark datasets in the literature Liu et al. (2023). This work contributes a new approach to spatial-temporal forecasting, potentially providing more accurate and reliable predictions for real-world applications such as traffic flow prediction. In future work, we plan to investigate the potential of applying time series foundation models Yeh et al. (2023a) for tackling spatial-temporal forecasting problems.
@@ Appendix A Supplementary
A.1 Complex Linear Layer
As depicted in Fig. 10, the time series in the large spatial-temporal forecasting benchmark dataset exhibits periodicity.
Figure 10: The time series for three randomly selected nodes from the LargeST dataset spans a period of four weeks.
Figure 11: The complex linear layer.
To leverage this periodicity, we opt to process the time series in the frequency domain using a complex linear layer Trabelsi et al. (2018). The computational graph for the complex linear layer is illustrated in Fig. 11. Initially, the input data is converted to the frequency domain using the Fast Fourier Transform (FFT) method. Subsequently, the real and imaginary parts are processed by two distinct linear layers, one containing the real part of the model weights and the other containing the imaginary part. The outputs from these two linear layers are then combined, and finally an inverse FFT layer is used to convert back to the time domain.
$(W_{\text{real}}x_{\text{real}}-W_{\text{imag}}x_{\text{imag}})+{i\mkern 1.0mu}(W_{\text{real}}x_{\text{imag}}+W_{\text{imag}}x_{\text{real}})$ (6)
This operation is captured in the design illustrated in Fig. 11.
A.2 Dataset Statistics
@@
Table 2: Dataset statistics.
@@ A.3 Baseline Method
  • LSTM Hochreiter and Schmidhuber (1997): The Long Short-Term Memory (LSTM), a variant of the recurrent neural network architecture, is specifically designed to process sequential data, making it a general method for time series forecasting. When deploying this architecture on the LargeST dataset, the same model weights are utilized across the time series of different nodes.
@@
  • ASTGCN Guo et al. (2019): The Attention-based Spatial-Temporal Graph Convolutional Network (ASTGCN) model employs a spatial-temporal attention mechanism to capture spatial-temporal correlations. It also uses graph convolutions and conventional convolutions to extract spatial and temporal patterns, respectively.
@@
  • DCRNN Li et al. (2017): The Diffusion Convolutional Recurrent Neural Network (DCRNN) incorporates diffusion convolution, the sequence-to-sequence architecture, and the scheduled sampling technique to capture spatial-temporal patterns.
@@
  • AGCRN Bai et al. (2020): The Adaptive Graph Convolutional Recurrent Network (AGCRN) extends the design of recurrent networks with a node adaptive parameter learning module, enabling the capture of node-specific patterns. Additionally, it includes a data adaptive graph generation module to infer the interdependencies among different nodes.
@@
  • STGCN Yu et al. (2017): The Spatial-Temporal Graph Convolutional Networks (STGCN) employs both graph and temporal convolutions to model the spatial-temporal correlations within the data.
@@
  • GWNET Wu et al. (2019): The Graph WaveNet (GWNET) is another network that relies on convolution operations. Specifically, it employs a gated temporal convolution module and a graph convolution layer to model the spatial-temporal correlation.
@@
  • STGODE Fang et al. (2021): The Spatial-Temporal Graph Ordinary Differential Equation Networks (STGODE) models spatial-temporal dynamics using tensor-based ordinary differential equations. The model design also incorporates a semantical adjacency matrix and temporal dilated convolution modules.
@@
  • DSTAGNN Lan et al. (2022): The Dynamic Spatial-Temporal Aware Graph Neural Network (DSTAGNN) is a method that does not use a predefined static adjacency matrix. Instead, it learns the dynamic spatial associations among nodes and utilizes a spatial-temporal attention module based on multi-order Chebyshev polynomials to capture these associations. To model the temporal associations, it employs gated convolution modules.
@@
  • DGCRN Li et al. (2023): The Dynamic Graph Convolutional Recurrent Network (DGCRN) model combines graph convolution networks with recurrent networks. In this model, a dynamic adjacency matrix is progressively generated by a hyper-network in synchronization with the recurrent steps. This dynamic adjacency matrix, in conjunction with the predefined static adjacency matrix, is utilized to generate predictions.
[…]
@@
  • 1NN: The 1-Nearest-Neighbor (1NN) method serves as a simple baseline for a range of time series problems Bagnall et al. (2017); Dau et al. (2019). We have implemented this baseline by leveraging the matrix profile for its efficiency Yeh et al. (2016); Zimmerman et al. (2019). Notably, the version of the 1NN method benchmarked in this study is the most basic form, where each node is treated as independent from the others. While more advanced techniques from the literature, such as those presented in Yeh et al. (2017); Martínez et al. (2019); Yeh et al. (2022b), could potentially enhance the method, we plan to explore these techniques in our future work.
@@
  • Linear: Having proven its effectiveness in general time series forecasting Zeng et al. (2023), the linear model has been incorporated into our benchmark experiments. Similar to LSTM, the same model weights are used across the time series of various nodes.
@@
  • TSMixer Chen et al. (2023): The Time Series Mixer (TSMixer) is a stacked Multi-Layer Perceptron (MLP) that efficiently extracts information by utilizing mixing operations across both the time and feature dimensions (i.e., nodes for spatial-temporal data). These mixing operations are capable of capturing the relationships between different time steps and nodes.
A.4 Implementation Detail
@@
The experiments were carried out using Python 3.10.11 on a Linux server equipped with an AMD EPYC 7713 64-Core Processor and NVIDIA Tesla A100 GPU. Our implementation leverages PyTorch 2.0.1 to realize the proposed RPMixer, TSMixer, and the Linear model. All three models are trained using the AdamW optimizer Loshchilov and Hutter (2017) with default hyper-parameter settings. The MAE loss function was used, following Liu et al. (2023). In the case of the proposed method, we configure the number of mixer blocks to eight and set the random projection dimension to $\sqrt{n}$, where $n$ is the number of nodes in the graph. For TSMixer, the number of mixer blocks is also set to eight, and the number of hidden dimensions is fixed at 64, in accordance with the parameterization used by the original author in the Traffic prediction dataset experiment Chen et al. (2023). The 1NN method is implemented using pyscamp 0.4.0 Zimmerman et al. (2019). Further details on our implementation can be found in the released source code The Author(2023) (s).
(a) RMSE
(b) MAPE
Figure 12: The correlation-error diagram illustrates the interplay of diversity and error (RMSE and MAPE) across individual blocks in the model.
A.5 Ablation Study

### v4

@@ 5.2. Baseline Method
5.3. Benchmark Result
The benchmark results are summarized in Table 1. Our discussion on the performance differences among various methods will proceed in three parts: firstly, we will discuss the 11 baselines included in the original LargeST benchmark (Liu et al., 2023b); next, we will address the three new time series forecasting baselines we added to the benchmark; finally, we will highlight the performance of our proposed RPMixer method in comparison to the baselines.
Regarding the original 11 baseline methods, HL and LSTM are typically outperformed by other methods as they do not consider inter-node relationships like the remaining nine spatial-temporal methods. Newer methods such as DGCRN and D2STGNN usually outperform other baselines. However, these two methods cannot be scaled to the two larger datasets, GLA and CA. Among older methods, AGCRN and GWNET outperform others on the SD, GBA, and GLA datasets. When considering the CA dataset, only four spatial-temporal models are applicable; GWNET, STGCN, and STGODE performance are comparable with each other.
In the case of the three new baselines (1NN, Linear, and TSMixer), both 1NN and Linear perform poorly for the same reason as the HL and LSTM baselines: these methods do not model the relationship between the nodes. When comparing TSMixer with the original 11 baseline methods, it is comparable with the best method on the two smaller datasets (SD and GBA). However, it performs noticeably worse in terms of RMSE on the two larger datasets (GLA and CA). A possible reason is that TSMixer is overfitting to MAE, which also serves as the loss function when training the model. Despite TSMixer’s high parameter count, it can still be applied to the largest dataset (CA) because the memory scales linearly with respect to the number of nodes, unlike the more expensive graph-based baselines (ASTGCN, AGCRN, DSTAGNN, DGCRN, and D2STGNN).
@@ 5.4. Ablation Study
Figure 9. The ablation study result measured using average MAE, RMSE, and MAPE. Removing each component resulted in a degradation of the performance.
The pre-activation design emerged as the most consequential design decision. Upon its replacement with a post-activation design, we effectively removed the most important element of the original design, leading to a significant performance decrease. This decline is likely due to the elimination of identity mapping connections from the network. These connections are vital for the ensemble interpretation of the residual connection. Without them, the inclusion of a random projection layer no longer makes sense, as it weakens the ensemble-like behavior of the residual blocks (i.e., the mixer blocks in RPMixer), which was the motivation behind the design of the random projection layer.
The random projection layer design was the second most important decision in terms of performance difference. We disabled the random projection layer by converting it into a regular linear layer with trainable weights. This layer helps each mixer block in focusing on different aspects of the inter-node relationship, so its importance is not surprising. To verify our claim that the random projection layer promotes diversity, we constructed a correlation-error diagram comparing the proposed method with the variant without random projection layer. This diagram serves a similar purpose to the kappa-error diagram (Margineantu and Dietterich, 1997) used for analyzing base learners in an ensemble model. Before discussing the diagram, let us first introduce what a correlation-error diagram is.
A correlation-error diagram is a visualization tool akin to the kappa-error diagram (Margineantu and Dietterich, 1997), used for analyzing the trade-off between performance and diversity in ensemble models. Each dot in the diagram represents a pair of base learners, with the $y$-axis showing the average performance of the pair’s outputs and the $x$-axis indicating the degree of agreement between them. Specifically, we use the Pearson correlation coefficient to measure agreement between the outputs of a pair of mixer blocks (i.e., the base learners). We use MAE, RMSE, and MAPE as performance measures. The figure created using MAE and Pearson correlation coefficient is shown in Fig. 10. Figures using other performance measures are included in Appendix A.5. We observe that the dots associated with the proposed method with a random projection layer have a wider distribution (larger range in $y$-axis) compared to the variant without the random projection layer. This confirms that our random projection layer effectively increases the diversity of the intermediate representation.
Figure 10. The correlation-error diagram illustrates the interplay of diversity and error (MAE) across individual blocks in the model.
@@
5.5. Parameter Sensitivity Analysis
In our parameter sensitivity analysis, we focus on two hyper-parameters: the number of mixer blocks and the number of neurons in the random projection layer. We report performance on both the validation and test sets to demonstrate the generalizability of our findings across different data partitions. Fig. 11 illustrates the parameter sensitivity analysis for the number of mixer blocks.
Figure 11. The parameter sensitivity analysis for the number of block parameter measured using average MAE, RMSE, and MAPE. Setting the number of block to eight generally yield better result. This observation holds true for both validation and test data.
@@ 5.5. Parameter Sensitivity Analysis
For the number of neurons in the random projection layer, we set it as a function with respect to the number of nodes in the spatial-temporal data, as more neurons may be required for datasets with larger number of nodes. If the graph has $n$ nodes, we set the number of neurons in the random projection layer as $m_{neuron}\sqrt{n}$, where $m_{neuron}$ is the hyper-parameter controlling the number of neurons in the layer. This design allows the model to automatically use more neurons for datasets with more number of nodes, even if the hyper-parameter $m_{neuron}$ is set to the same value for all datasets. Fig. 12 displays the results of the parameter sensitivity analysis for the hyper-parameter $m_{neuron}$.
Figure 12. The parameter sensitivity analysis for the number of neurons parameter measured using average MAE, RMSE, and MAPE. The number is set by $m_{neuron}\sqrt{n}$, where $n$ is the number of nodes in the graph and $m_{neuron}$ is a factor. Setting this factor to one often has the best performance. This observation holds true for both validation and test data.
@@
6. Conclusion
In this paper, we proposed RPMixer, an all-MLP mixer model that incorporates random projection layers. These layers enhance the diversity among the outputs of each mixer block, thereby improving the overall performance of the model. Our experiments demonstrated that the random projection layers not only improve the diversity of the intermediate representation but also boost the model’s overall performance on large-scale spatial-temporal benchmark datasets in the literature (Liu et al., 2023b). In future work, we plan to investigate the potential of applying time series foundation models (Yeh et al., 2023a) for tackling spatial-temporal forecasting problems.
@@ Appendix A Supplementary
A.1. Dataset Statistics
@@
Table 2. Dataset statistics.
@@ A.2. Baseline Method
  • LSTM (Hochreiter and Schmidhuber, 1997): The Long Short-Term Memory (LSTM), a variant of the recurrent neural network architecture, is specifically designed to process sequential data, making it a general method for time series forecasting. When deploying this architecture on the LargeST dataset, the same model weights are utilized across the time series of different nodes.
@@
  • ASTGCN (Guo et al., 2019): The Attention-based Spatial-Temporal Graph Convolutional Network (ASTGCN) model employs a spatial-temporal attention mechanism to capture spatial-temporal correlations. It also uses graph convolutions and conventional convolutions to extract spatial and temporal patterns, respectively.
@@
  • DCRNN (Li et al., 2017): The Diffusion Convolutional Recurrent Neural Network (DCRNN) incorporates diffusion convolution, the sequence-to-sequence architecture, and the scheduled sampling technique to capture spatial-temporal patterns.
@@
  • AGCRN (Bai et al., 2020): The Adaptive Graph Convolutional Recurrent Network (AGCRN) extends the design of recurrent networks with a node adaptive parameter learning module, enabling the capture of node-specific patterns. Additionally, it includes a data adaptive graph generation module to infer the inter-dependencies among different nodes.
@@
  • STGCN (Yu et al., 2017): The Spatial-Temporal Graph Convolutional Networks (STGCN) employs both graph and temporal convolutions to model the spatial-temporal correlations within the data.
@@
  • GWNET (Wu et al., 2019): The Graph WaveNet (GWNET) is another network that relies on convolution operations. Specifically, it employs a gated temporal convolution module and a graph convolution layer to model the spatial-temporal correlation.
@@
  • STGODE (Fang et al., 2021): The Spatial-Temporal Graph Ordinary Differential Equation Networks (STGODE) models spatial-temporal dynamics using tensor-based ordinary differential equations. The model design also incorporates a semantical adjacency matrix and temporal dilated convolution modules.
@@
  • DSTAGNN (Lan et al., 2022): The Dynamic Spatial-Temporal Aware Graph Neural Network (DSTAGNN) is a method that does not use a predefined static adjacency matrix. Instead, it learns the dynamic spatial associations among nodes and utilizes a spatial-temporal attention module based on multi-order Chebyshev polynomials to capture these associations. To model the temporal associations, it employs gated convolution modules.
@@
  • DGCRN (Li et al., 2023a): The Dynamic Graph Convolutional Recurrent Network (DGCRN) model combines graph convolution networks with recurrent networks. In this model, a dynamic adjacency matrix is progressively generated by a hyper-network in synchronization with the recurrent steps. This dynamic adjacency matrix, in conjunction with the predefined static adjacency matrix, is utilized to generate predictions.
  • D2STGNN (Shao et al., 2022d): The Decoupled Dynamic Spatial-Temporal Graph Neural Network (D2STGNN) decouples the diffusion and inherent information in a data-driven manner using an estimation gate and a residual decomposition mechanism. Additionally, it employs a dynamic graph learning module that learns the dynamic characteristics of the spatial-temporal graph.
@@
  • 1NN: The 1-Nearest-Neighbor (1NN) method serves as a simple baseline for a range of time series problems (Bagnall et al., 2017; Dau et al., 2019). We have implemented this baseline by leveraging the matrix profile for its efficiency (Yeh et al., 2016; Zimmerman et al., 2019). Notably, the version of the 1NN method benchmarked in this study is the most basic form, where each node is treated as independent from the others. While more advanced techniques from the literature, such as those presented in (Yeh et al., 2017; Martínez et al., 2019; Yeh et al., 2022b), could potentially enhance the method, we plan to explore these techniques in our future work.
@@
  • Linear: Having proven its effectiveness in general time series forecasting (Zeng et al., 2023), the linear model has been incorporated into our benchmark experiments. Similar to LSTM, the same model weights are used across the time series of various nodes.
@@
  • TSMixer (Chen et al., 2023): The Time Series Mixer (TSMixer) is a stacked Multi-Layer Perceptron (MLP) that efficiently extracts information by utilizing mixing operations across both the time and feature dimensions (i.e., nodes for spatial-temporal data). These mixing operations are capable of capturing the relationships between different time steps and nodes.
A.3. Implementation Detail
@@
The experiments were carried out using Python 3.10.11 on a Linux server equipped with an AMD EPYC 7713 64-Core Processor and NVIDIA Tesla A100 GPU. Our implementation leverages PyTorch 2.0.1 to realize the proposed RPMixer, TSMixer, and the Linear model. All three models are trained using the AdamW optimizer (Loshchilov and Hutter, 2017) with default hyper-parameter settings. The MAE loss function was used, following (Liu et al., 2023b). We have incorporated an early stopping mechanism, with the patience parameter set to seven. In the case of the proposed method, we configure the number of mixer blocks to eight and set the random projection dimension to $\sqrt{n}$, where $n$ is the number of nodes in the graph. For TSMixer, the number of mixer blocks is also set to eight, and the number of hidden dimensions is fixed at 64, in accordance with the parameterization used by the original author in the Traffic prediction dataset experiment (Chen et al., 2023). The 1NN method is implemented using pyscamp 0.4.0 (Zimmerman et al., 2019). 1NN was not tuned because there are no hyper-parameters in our implementation. Further details on our implementation can be found in the released source code (The Author(2023), s). The results of LSTM (Hochreiter and Schmidhuber, 1997), ASTGCN (Guo et al., 2019), AGCRN (Bai et al., 2020), DSTAGNN (Lan et al., 2022), DGCRN (Li et al., 2023a), and D2STGNN (Shao et al., 2022d) were obtained from (Liu et al., 2023b).
A.4. Additional Discussion about Table 1
In this section, we delve into a more comprehensive discussion about the benchmark results shown in Table 1.
Fist, we examine the different graph neural network-based methods. This analysis is performed using the average performances. The best graph method on the SD dataset is D2STGNN, which outperforms the others on both MAE and RMSE. For MAPE, DGCRN outperforms the others. Both D2STGNN and DGCRN have an adjacency matrix learning component, suggesting that modeling the cross-variate dependency is important for the forecasting problem. For the GBA dataset, AGCRN, GWNET, and D2STGNN achieve the best MAE, RMSE, and MAPE, respectively. These methods once again consist of an adjacency matrix learning component. For the GLA dataset, AGCRN is the best graph method, achieving the best MAE and MAPE, while GWNET achieves the best RMSE. Both methods are capable of adjacency matrix learning. For the CA dataset, STGODE, DCRNN, and STGCN achieve the best MAE, RMSE, and MAPE, respectively. These three methods outperform the others because the stronger graph alternatives are not applicable for the CA dataset due to scalability issues.
Next, we compared the proposed RPMixer with the second-best results on different datasets. According to the average MAE, RMSE, and MAPE, the second-best results on the SD dataset are from D2STGNN and DGCRN, the second-best results on the GBA dataset are from D2STGNN and TSMixer, the second-best results on the GLA dataset are from AGCRN and GWNET, and the second-best results on the CA dataset are from GWNET and TSMixer. We are providing a detailed analysis of the similarities and differences between our proposed method and AGCRN, D2STGNN, DGCRN, GWNET, and TSMixer in the following.
AGCRN is a graph convolutional recurrent network with two special designs: 1) node adaptive parameter learning and 2) data adaptive graph generation. The major difference between AGCRN and our method is that AGCRN uses graph convolution and recurrent operations to process the input signal, while our method uses only linear layers to process the signal. AGCRN utilizes data adaptive graph generation techniques to learn the dependency between different nodes, while our method learns such information with spatial mixers. However, for both methods, the ability to learn inter-node dependency is crucial for achieving competitive performances. Note that the data adaptive graph generation module needs to compute an $n$-by-$n$ node similarity matrix; therefore, the major advantage of our proposed method compared to AGCRN is that our method has a smaller memory cost as it does not compute any $n$-by-$n$ matrices.
D2STGNN decouples the signal into two parts with an estimation gate. The first part is processed with diffusion convolutional layers and the second part is processed with gated recurrent units (GRU) layers. In addition, D2STGNN has a dynamic graph learning model which learns the $n$-by-$n$ dynamic transition matrices that indicate the relationship between different nodes. One similarity between D2STGNN and our proposed method is that D2STGNN also has an ensemble-like structure where the decoupled signals are processed with two different models (i.e., the convolutional layers and GRU layers). This suggests that an ensemble-like structure could be important for achieving competitive performances on spatial-temporal datasets. D2STGNN learns the dependency between different nodes with the diffusion convolutional layers, and our method learns such information with spatial mixers. The ability to learn inter-node dependency is crucial for achieving competitive performances. The major advantage of our proposed method compared to D2STGNN is that our method has a smaller memory cost as it does not compute any $n$-by-$n$ matrices.
DGCRN is a graph convolutional recurrent network capable of generating dynamic graphs (i.e., $n$-by-$n$ dynamic adjacency matrices to capture inter-node dependency. Our method internally learns the inter-node dependency with spatial mixers. For both methods, the ability to learn inter-node dependency is crucial for achieving competitive performances. The major difference between DGCRN and our method is that DGCRN uses graph convolution and recurrent operations to process the input signal, while our method uses only linear layers to process the signal. The major advantage of our proposed method compared to DGCRN is that our method has a smaller memory cost as it does not compute any $n$-by-$n$ matrices.
GWNET consists of several convolutional blocks, each containing two types of layers: 1) gated temporal convolutional layers and 2) diffusion convolutional layers. The major difference between the two methods is that GWNET uses convolutional layers to process the signals, whereas our proposed method uses linear layers. The first similarity shared by GWNET and our proposed method is that both methods adopt residual layers. In other words, the blocks in GWNET could have ensemble-like behavior, and such design may contribute greatly to the success of both methods. Another similarity is that D2STGNN learns the dependency between different nodes with the diffusion convolutional layers, and our method learns such information with spatial mixers. The ability to learn inter-node dependency is also crucial for achieving competitive performances. Because GWNET computes $n$-by-$n$ self-adaptive adjacency matrix in diffusion convolutional layers, the major advantage of our proposed method compared to GWNET is that our method has a smaller memory cost as it does not compute any $n$-by-$n$ matrices.
TSMixer is the method most similar to ours. The two differences are the complex linear layers and random projection layers. According to the ablation study (see Section 5.4), the majority of the performance gain came from the random projection layers.
In conclusion, the capability of learning inter-node dependency is crucial for the success of a method. The random projection layer design is the major reason for our model to outperform the baseline method (i.e., TSMixer) that is most similar to ours.
A.5. Ablation Study